OpenStudy (anonymous):

Consider the equation 7x + 3y = 42

OpenStudy (anonymous):

On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points)

OpenStudy (anonymous):

u r working it out

OpenStudy (richyw):

solve for y\[3y=42-7x\]\[y=\frac{42}{3}-\frac{7}{3}x\]putting this in the familiar y=mx+b:\[y=-\frac{7}{3}x+\frac{42}{3}=\frac{7}{3}x+14\]To graph a straight line, you only need two points, or one point and a slope. you already have the slope (-7/3) and a point in the last equation, but I will still just use the two points method. The easiest points to choose are the x and y intercepts. For the y intercept, set \(x=0\) \[y=\frac{-7}{3}(0)+14\]\[y=14\] for the x intersect, set \(y=0\) and solve\[0=\frac{-7}{3}x+14\]\[\frac{7}{3}x=14\]\[\frac{1}{3}x=\frac{14}{7}=2\]\[x=2*3=6\] so you now have two points on your graph \((0,14), (7,0)\) now you must simply plot those two points and connect the lines

OpenStudy (richyw):

ok sorry I didn't see there were two methods. I kinda explained that in the first paragraph. The method I showed after that was the "intercepts method" for the point-slop method you just need to find a point (one of those intercepts will work fine) and then just draw a line through it with the correct slope \(\frac{-7}{3}\)

OpenStudy (richyw):

remember that slope is \[\frac{rise}{run}\] so for every 7 that you go along the x axis in the postive direction, the you will go down 3 along the y axis. so just plot a point that is 7 to the right and three below of the other point you have chosen, and connect the two lines. I must go now as I have an exam, but hope that helps.